If an air number density is n=10^19cm^-³ and T=300K.Justify weather this Air is plasma or not?please solve it completely......


1
Expert's answer
2020-07-30T10:50:31-0400

Consider one cubic centimetre of air with a molar mass of 29 g/mol and density at 300 K of 1.18 g/cm3. Therefore, the number of moles will be


"n=\frac{m}{\Mu}=\frac{\rho V}{\Mu}=\frac{1.18\cdot1}{1}=1.18\text{ mol}."

The number of particles in this one cubic centimetre will be


"N=nN_A=1.18\cdot6.022\cdot10^{23}=\\=7.106\cdot10^{23}\text{ cm}^{-3}."

Therefore, since the substance in the condition of the problem contains 10 000 times fewer particles than air, this is plasma.